# Automated threshold selection and associated inference uncertainty for univariate extremes

### v4

@@ 1 Introduction
An inherent challenge in risk modelling is the estimation of high quantiles, known as return levels, beyond observed values. Such inference is important for designing policies or protections against future extreme events, e.g., in finance or hydrology (Smith,, 2003; Coles et al.,, 2003). Extreme value methods achieve this extrapolation by using asymptotically exact models to approximate the tail of a distribution above a high, within-sample, threshold. They require the selection of a threshold $u$, such that above $u$, extreme value methods provide a suitable model. Here, we develop novel methods for automatic selection of the threshold and for propagating the uncertainty in this selection into return level inferences.
@@ 2 Background
For the Case 4 data, the plot shows that candidate thresholds above (below) $0.3$ are possibly appropriate (not appropriate) as CIs for higher candidate thresholds include (exclude) the corresponding shape parameter estimates, and above 0.8 the point estimates appear more stable. Here $(u,\xi)=(1,0.1)$, so we can see that candidates below 0.3 are not suitable as $\xi$ is outside their CIs, but the true threshold is higher than may be selected using this plot. For the River Nidd, lower candidate threshold values imply a very heavy-tailed distribution ($\hat{\xi}\approx 0.5$), whilst high candidate thresholds imply a very short tail, with estimates ($\hat{\xi}\approx-1$) indicating that the distribution’s upper end point is essentially the largest observation. As a result of this unusual behaviour, the Nidd data has become the primary example for non-trivial threshold selection (Davison and Smith,, 1990; Northrop and Coleman,, 2014). We apply our new method to this dataset in Section 7. Further examples of the problems encountered when using parameter stability plots are given in supplementary material S:2.2.
@@ 4.2 Investigation of the EQD metric choice
We choose to define $(p_{j},q_{i})$ as the standard plotting points for a QQ-plot assessment and to weight contributions to $d_{b}(u)$ equally across the quantiles. Although higher (lower) sample quantiles exhibit greater (less) sampling variability, equal weighting is appropriate when taking the $\{p_{j}\}$ values to be equally-spaced because for any $\xi>-1$, the GPD density is monotonically decreasing. This leads to dense evaluation for lower sample quantiles and more sparse evaluation in the upper tail. The choice of equal weighting on this scale is motivated and supported by empirical evidence in Varty et al., (2021). We considered other $(p_{j},q_{i})$ definitions but found that there was no systematic ordering of the performance over these definitions and any differences in RMSE for the thresholds selected by the EQD were minimal, especially when compared to the differences between the EQD and existing methods.
@@ 5 Accounting for parameter and threshold uncertainty
GPD inferences are sensitive to the choice of threshold (Davison and Smith,, 1990) but uncertainty about this choice is not represented in Algorithms 1 or 1b. This omission would be important when return levels inform the design of hazard protection mechanisms, where omitting this source of uncertainty could lead to over-confidence in the inference and have dangerous consequences. Algorithm 2 provides a novel method to propagate both threshold and parameter uncertainty through to return level estimation, using a double-bootstrap procedure. To focus on the threshold uncertainty and to forgo the need for a parametric model below the threshold, we employ a non-parametric bootstrap procedure on the original dataset. We resample with replacement $n$ values from the observed data $B_{2}$ times, estimate a threshold for each such bootstrap sample using the automated selection method of Section 4, and fit a GPD to the excesses of this threshold. For each one of the $B_{2}$ samples, we employ Algorithm 1 to account for the subsequent uncertainty in the GPD parameters. Calculating a summary statistic for each of the $B_{1}\times B_{2}$ samples leads to a distribution of bootstrapped estimates that accounts for uncertainty in the threshold selection as well as in the GPD and threshold exceedance rate parameters. We use $B_{1}=B_{2}=200$. In Section 6, we illustrate how using Algorithm 2 improves the coverage probability of CIs, and in Section 7 how it widens CIs for return levels of the River Nidd.
@@ 6.2 Scenario 1: True threshold for GPD tail
Cases 1-3 all have a distinct changepoint in the density at the true threshold. Cases 1 and 2 have the same distribution, with $\xi>0$, with Case 2 having a smaller sample size. We find that the Wadsworth method fails to estimate a threshold in samples with $\xi<-0.05$ irrespective of sample size, so Case 3 is selected near that boundary where the method works reasonably and has double the sample size of Case 1. Case 4 provides a more difficult example with a continuous density and a small number of exceedances of the true threshold. The data are derived from a partially observed GPD, denoted GPDp, with data from a GPD above 0 and rejected if less than an independent realisation from a Beta(1,2) distribution.
@@
Cases 1-4, Threshold recovery: Table 2 shows the RMSE of the chosen thresholds for each method in Cases 1-4. The EQD achieves RMSEs 1.2-7.7 (1-11.2) times smaller than the Wadsworth (Northrop) method. The EQD has the lowest bias by a considerable margin in Cases 1-3 and shows the lowest variance in threshold estimation in all cases. In fact, the variance is reduced by factor of at least 20 relative to both the Wadsworth and Northrop methods (see Table S.9).
@@
Cases 1-4, Quantile recovery: Table 3 presents the RMSEs for the $(1-p_{j,n})$-quantiles where $p_{j,n}=1/(10^{j}n)$ for $j=0,1,2$ for sample size $n$, which ensures that extrapolation is equally difficult over $n$ for a given $j$. When $j=0$, no extrapolation is required so the choice of $u$ should not be too important; the similar RMSEs across methods reflect this. As $j$ increases, all RMSEs increase and the differences between methods become clear. The EQD method is best uniformly, followed by the Wadsworth and then the Northrop methods. This pattern reflects the findings in Table 2, although here with differential performances sensitive to $j$. The EQD achieves the lowest bias in the majority of cases and leads to quantile estimates with considerably less variance in all cases, particularly as $j$ increases.
@@ 7 Application to river flow data
Comparing thresholds selected between the methods is complicated due to the sensitivity of the Wadsworth and Northrop methods to the grid choice. For the EQD, it is natural to use the densest and widest grid, giving $\hat{u}=67.10$. This threshold, which is lower than previously found, gives far more data for the extreme value analysis. As all the River Nidd data are “extreme”, we believe taking $u$ so close to the lower endpoint of the data is not problematic, and it may indicate that the pre-processing level used to produce these data was too high. The first estimated threshold from the Wadsworth (Northrop) methods, without convergence or warning issues, is $\hat{u}=149.10$ ($\hat{u}=65.08$). For these three threshold choices, the corresponding GPD parameter estimates (and 95% CIs) are: $\hat{\sigma}_{u:EQD}=23.74~{}(17.78,29.70)$ and $\hat{\xi}=0.26~{}(0.06,0.46)$ for the EQD; $\hat{\xi}=-0.15~{}(-1.00,0.70)$ for the Wadsworth method; and for the Northrop method, $\hat{\xi}=0.20~{}(0.02,0.38)$, where we omit the latter two scale parameter estimates as they are estimating different quantities which depend on the threshold, see Section 2. Provided all estimated thresholds are high enough for the GPD to be appropriate, the values of $\hat{\xi}$ should be similar across methods, due to the threshold stability property (see supplementary material S:2). The Wadsworth method has an extremely wide CI, which leads to meaningless inference. However, the EQD and Northrop findings about $\xi$ are very similar, but the sensitivity to the candidate grid is still a problem for the Northrop method.

### v5

@@ 1 Introduction
An inherent challenge in risk modelling is the estimation of high quantiles, known as return levels, beyond observed values. Such inference is important for designing policies or protections against future extreme events, e.g., in finance or hydrology (Smith,, 2003; Coles et al.,, 2003). Extreme value methods achieve this extrapolation by using asymptotically exact models to approximate the tail of a distribution above a high, within-sample, threshold $u$. The choice of this threshold is fundamental in providing meaningful inference. Here, we develop novel methods for automatic selection of the threshold and for propagating the uncertainty in this selection into return level inferences.
@@ 2 Background
For the Case 4 data, the plot shows that candidate thresholds above (below) $0.3$ are possibly appropriate (not appropriate) as CIs for higher candidate thresholds include (exclude) the corresponding shape parameter estimates, and above 0.8 the point estimates appear more stable. Here $(u,\xi)=(1,0.1)$, so we can see that candidates below 0.3 are not suitable as $\xi$ is outside their CIs, but the true threshold is higher than may be selected using this plot. For the River Nidd, lower candidate threshold values imply a very heavy-tailed distribution ($\hat{\xi}\approx 0.5$), whilst high candidate thresholds imply a very short tail, with estimates ($\hat{\xi}\approx-0.5$). As a result of this unusual behaviour, the Nidd data has become the primary example for non-trivial threshold selection (Davison and Smith,, 1990; Northrop and Coleman,, 2014). We apply our new method to this dataset in Section 7. Further examples of the problems encountered when using parameter stability plots are given in supplementary material S:2.2.
@@ 4.2 Investigation of the EQD metric choice
We choose to take $\{p_{j}\}$ to be equally-spaced and to weight contributions to $d_{b}(u)$ equally across the corresponding quantiles. Although higher (lower) sample quantiles exhibit greater (less) sampling variability, equal weighting is appropriate when taking the $\{p_{j}\}$ values to be equally-spaced because for any $\xi>-1$, the GPD density is monotonically decreasing. This leads to dense evaluation for lower sample quantiles and more sparse evaluation in the upper tail. The choice of equal weighting on this scale is motivated and supported by empirical evidence in Varty et al., (2021). Our choice for $p_{j}$ is based on the expression for plotting points in a QQ-plot assessment we use in previous research and the choice for $q_{i}$ is the default option for the R quantile function. As these choices are subjective, we also consider alternative definitions but find that there is no systematic ordering of the performance over these definitions and any differences in RMSE for the thresholds selected by the EQD are minimal, especially when compared to the differences between the EQD and existing methods.
@@ 5 Accounting for parameter and threshold uncertainty
GPD inferences are sensitive to the choice of threshold (Davison and Smith,, 1990) but uncertainty about this choice is not represented in Algorithms 1 or 1b. This omission would be important when return levels inform the design of hazard protection mechanisms, where omitting this source of uncertainty could lead to over-confidence in the inference and have dangerous consequences. Algorithm 2 provides a novel method to propagate both threshold and parameter uncertainty through to return level estimation, using a double-bootstrap procedure. To focus on the threshold uncertainty and to forgo the need for a parametric model below the threshold, we employ a non-parametric bootstrap procedure on the original dataset. We resample with replacement $n$ values from the observed data $B_{2}$ times, estimate a threshold for each such bootstrap sample using the automated selection method of Section 4, and fit a GPD to the excesses of this threshold. For each one of the $B_{2}$ samples, we employ Algorithm 1 to account for the subsequent uncertainty in the GPD parameters. Calculating a summary statistic for each of the $B_{1}\times B_{2}$ samples leads to a distribution of bootstrapped estimates that accounts for uncertainty in the threshold selection as well as in the GPD and threshold exceedance rate parameters. We use $B_{1}=B_{2}=200$. To run this algorithm using the EQD method for the threshold selection step (which itself has $B$ bootstraps), it would require $B_{2}(B+B_{1})$ bootstrap samples to be generated. Specifically, for the $B_{2}$ samples initially generated for Algorithm 2, we have $B_{2}\times B$ in selecting the threshold values and $B_{2}\times B_{1}$ in capturing the GPD parameter uncertainty above these selected thresholds. this can res In Section 6, we illustrate how using Algorithm 2 improves the coverage probability of CIs, and in Section 7 how it widens CIs for return levels of the River Nidd.
@@ 6.2 Scenario 1: True threshold for GPD tail
Cases 1-3 all have a distinct changepoint in the density and density mode both at the true threshold which should make all methods of threshold selection perform better than in situations without either of these features. Cases 1 and 2 have the same distribution, with $\xi>0$, with Case 2 having a smaller sample size. We find that the Wadsworth method fails to estimate a threshold in samples with $\xi<-0.05$ irrespective of sample size, so Case 3 is selected near that boundary where the method works reasonably and has double the sample size of Case 1. Case 4 provides a more difficult example with a continuous density and a small number of exceedances of the true threshold. The data are derived from a partially observed GPD, denoted GPDp, with data drawn from a GPD above 0 and rejected if less than an independent realisation from a Beta(1,2) distribution.
@@
Cases 1-4, Threshold recovery: Table 2 shows the RMSE of the chosen thresholds for each method in Cases 1-4, with the EQD achieving RMSEs 1.2-7.7 (1-11.2) times smaller than the Wadsworth (Northrop) method. The EQD has the lowest bias by a considerable margin in Cases 1-3 and shows the lowest variance in threshold estimation in all cases. In fact, the variance is reduced by a factor of at least 20 relative to both the Wadsworth and Northrop methods (see Table S.9). The very strong performance of the EQD relative to both the Wadsworth and Northrop methods is particularly noteworthy in Cases 1-3, and is also seen for Case 0 and later for Cases 5-7. We believe that the key reason for this is the discontinuity in the density, a feature common to all of these cases, as that appears to lead to a very small bias for the EQD method relative to the other methods. Specifically, the variance penalty of the EQD metric seems to push the threshold as low as possible, but its complementary goodness-of-fit measure almost entirely stops the threshold being selected below the clear discontinuity in the density. For Case 4, which has a continuous density, the EQD achieves the smallest RMSE almost entirely due to it having the smallest variance but with a bias component broadly comparable with the other methods.
@@
Cases 1-4, Quantile recovery: Table 3 presents the RMSEs for the $(1-p_{j,n})$-quantiles where $p_{j,n}=1/(10^{j}n)$ for $j=0,1,2$ for sample size $n$, which ensures that extrapolation is equally difficult over $n$ for a given $j$. When $j=0$, no extrapolation is required so the choice of $u$ should not be too important; the similar RMSEs across methods reflect this. As $j$ increases, all RMSEs increase and the differences between methods become clear. The EQD method is best uniformly, followed by the Wadsworth and then the Northrop methods. This pattern reflects the findings in Table 2, although here with differential performances sensitive to $j$. However, in terms of quantile estimation, the EQD method does not retain the large differential relative to the other methods which was seen for threshold selection in Cases 1-3. In contrast, the differences between methods in Case 4 are now more apparent, as controlling the variance is more important than any small differences in bias when we are concerned with a RMSE assessment of quantiles which lie far into the tail. The EQD achieves the lowest bias in the majority of cases and leads to quantile estimates with considerably less variance in all cases, particularly as $j$ increases.
@@ 7 Application to river flow data
Comparing thresholds selected between the methods is complicated due to the sensitivity of the Wadsworth and Northrop methods to the grid choice. For the EQD, it is natural to use the densest and widest grid, giving $\hat{u}=67.10$. This threshold, which is lower than previously found, gives far more data for the extreme value analysis. As all the River Nidd data are “extreme”, we believe taking $u$ so close to the lower endpoint of the data is not problematic, and it may indicate that the pre-processing level used to produce these data was too high. The first estimated threshold from the Wadsworth (Northrop) methods, without convergence or warning issues, is $\hat{u}=149.10$ ($\hat{u}=65.08$). For these three threshold choices, the corresponding GPD parameter estimates (and 95% CIs) are: $\hat{\sigma}_{u:EQD}=23.74~{}(17.78,29.70)$ and $\hat{\xi}=0.26~{}(0.06,0.46)$ for the EQD; $\hat{\xi}=-0.15~{}(-1.00,0.70)$ for the Wadsworth method; and for the Northrop method, $\hat{\xi}=0.20~{}(0.02,0.38)$, where we omit the latter two scale parameter estimates as they are estimating different quantities which depend on the threshold, see Section 2. Provided all estimated thresholds are high enough for the GPD to be appropriate, the values of $\hat{\xi}$ should be similar across methods, due to the threshold stability property (see supplementary material S:2). The Wadsworth method leads to an extremely wide CI, which results in meaningless inference. However, the EQD and Northrop findings about $\xi$ are very similar, but the sensitivity to the candidate grid is still a problem for the Northrop method.
@@ Acknowledgements
Conor Murphy, Jonathan A. Tawn
Department of Mathematics and Statistics, Lancaster University
and
Zak Varty
Department of Mathematics, Imperial College London